OpenStudy (anonymous):

A ball thrown upwards takes 4 s to reach the maximum height. Find (a) the initial speed with which it was thrown, and (b) the maximum height reached

OpenStudy (anonymous):

time taken to reach the maximum height is nothing but time of flight it is u/g..so we will get the value of initial velocity and maximum height is u^2/2g

OpenStudy (joannablackwelder):

I would find this using the formula Accelleration*time=change in velocity

OpenStudy (joannablackwelder):

Keep in mind that at the max height, the velocity will be 0. Make sense?

OpenStudy (anonymous):

yeah at max height final velocity will be zero..but we have the formula v^2-u^2 = 2gh since final velocity is zero substitute zero in place of v then we get -u^2=2gh but as the ball is thrown upwards take the sign of acceleration due to gravity as negative so u will get the max height as h =u^2/2g